OpenStudy (love.kat123):

Simplify the square root of 50. A. 5 times the square root of 2 B. the square root of 5 times the square root of 10 C. 2 times the square root of 5 D. 25

OpenStudy (love.kat123):

i think it might be D.

OpenStudy (anonymous):

\[\sqrt{50}\rightarrow \sqrt{25}*\sqrt{2}\rightarrow 5\sqrt{2}\]

OpenStudy (love.kat123):

so its A..............

OpenStudy (anonymous):

\[\sqrt{50}\approx7.07 \rightarrow 5\sqrt{2}\approx 7.07\]

OpenStudy (love.kat123):

okay thanks

OpenStudy (anonymous):

optionA..

OpenStudy (anonymous):

Np =0)